# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws (Extended Version)

### v2

@@ Abstract
Most existing necessary conditions for optimal control based on adjoining methods require both state and costate information, yet the unobservability of costates for a given feasible trajectory impedes the determination of optimality in practice. This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems with a single input, proposing the augmented switching law (ASL) that represents the input control and the feasibility in a compact form. Given a feasible trajectory, the perturbed trajectory under the constraints of ASL is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information. A first-order necessary condition is proposed that the Jacobian matrix of the ASL is not full row rank, which also results in a potential approach to optimizing a given feasible trajectory with the preservation of arc structures. The proposed necessary condition is applied to high-order chain-of-integrator systems with full box constraints, contributing to some theoretical results challenging to reason by costate-based conditions.
@@ I Introduction
Based on PMP, one can get rid of dependence on costates by deriving the optimal trajectory’s structure. Switching surfaces can be constructed in the state space for time-optimal control of linear systems. Walther et al. [17] computated switching surfaces for some systems through Gröebner basis. Patil et al. [18] represented switching surfaces using state equations for single-input diagonal linear systems without state constraints. However, for systems with state constraints, it remains challenging to completely construct switching surfaces due to the lack of optimal switching laws. For example, He et al. [19] provided explicit equations for switching surfaces of 3rd-order chain-of-integrator systems with full state constraints, limited to zero terminal states. Wang et al. [20] extended [19] to non-zero terminal conditions, failing to achieve time-optimality for higher-order systems. Numerical methods [21] face the curse of dimensionality. Hence, there is a lack of efficient and general optimality conditions without dependence on costates for high-order systems with state constraints.
Another type of necessary condition leverages Bellman’s principle of optimality [22], asserting that a sub-arc of an optimal trajectory is itself optimal for the induced sub-problem, without explicitly involving costates. The Hamilton-Jacobi-Bellman (HJB) equation [23] represents a typical necessary condition in the form of partial differential equations, which is widely used in optimal control [24, 25]. Evans and James [26] investigated the HJB equation for time-optimal control. Wolenski and Zhuang [27] showed that the minimal time function is the unique proximal solution to the HJB equation. However, most numerical HJB-based methods require high computational cost [28] due to the discretized grids. From a theoretical perspective, HJB-based methods also fail to preserve arc structures, which hinders the analytical reasoning.
Compared with the PMP-based and HJB-based necessary conditions, this paper sets out to establish a theoretical framework for time-optimal control of controllable single-input linear systems and develops a novel state-centric necessary condition that requires state information without dependence on costate information. Sections III and IV derive trajectories’ structures on arcs and keypoints for feasibility, respectively. Section V develops the state-centric necessary condition which is applied from theoretical and numerical perspectives in Sections VI and VII as examples, respectively. The contributions of this paper are as follows:
@@
    Based on the developed framework, this paper proposes a novel state-centric necessary condition for time-optimality, asserting that the Jacobian matrix induced by the ASL is not full row rank. The condition is derived from the fact that if a BBS trajectory is optimal, then all perturbed feasible trajectories have a strictly longer terminal time than the original one. Compared to existing costate-based conditions, the proposed state-centric necessary condition only requires state information without dependence on costates. Note that costates are usually unobservable in practice, especially when the trajectory is not planned by costate-based optimization.
@@
    This paper proposes a potential optimization idea based on the developed necessary condition. Given a feasible BBS trajectory as the initial value, the motion time is optimized with a fixed ASL until the Jacobian matrix is not full row rank. During iteration, the feasibility is guaranteed based on the developed ASL framework. Compared to discretized methods, the proposed approach can preserve the BBS arc structure and has the potential to significantly reduce control oscillations and open-loop integration errors of terminal states.
@@
    The proposed state-centric necessary condition is applied to high-order chain-of-integrator systems with full state constraints which remains an open problem in the optimal control domain. An upper bound on the number of arcs is determined in a general sense, while case-by-case analysis is necessary in traditional methods. A recursive equation of junction time is reasoned in chattering induced by 2nd-order constraints, which is challenging to prove by costate analysis. As a result, one could determine the existence of chattering induced by 2nd-order constraints in chain-of-integrator systems in future works. The above results demonstrate the theoretical effectiveness of the proposed necessary condition.
@@ II Problem Formulation
$\min_{u}\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t={t_{\mathrm{f}}},$ (1a)
@@
$u\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ is undetermined during ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, which is called a singular arc. The notation $\equiv$” means that the equality holds for a continuous period. (7) is the well-known bang-bang and singular (BBS) control law [11, 29]. In other words, the optimal control is always ±$\pm{u_{\mathrm{m}}}$, except during singular arcs.
Note that the objective function $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t$ is in a Lagrangian form; hence, the continuity of the system is guaranteed by the following equality, i.e.,
@@ III Arc Representation of the Optimal Trajectory
An arc of a trajectory is defined as a continuous segment conditioned by a fixed dynamic equation, i.e., $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ where $\widehat{{\boldsymbol{A}}}$ and $\widehat{{\boldsymbol{b}}}$ are constant. Two kinds of arcs, i.e., unconstrained arcs and constrained arcs, are discussed in Section III-A and Section III-B, respectively. Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-C. Finally, an optimal non-chattering trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-D.
III-A Unconstrained Arcs in Problem (1a)
In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds almost everywhere (a.e.), and the control does not switch. The following proposition states that $u\equiv\pm{u_{\mathrm{m}}}$ in an unconstrained arc.
@@ III-A Unconstrained Arcs in Problem (1a)
Proof.
Assume that a singular condition holds in the unconstrained arc, i.e., ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period. By (5), ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ implies that ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$ holds a.e. in an unconstrained arc. Hence, (4) implies that $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$. Therefore,
@@
By $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$, either ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ for a period or ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0 for finite times, while the former case is ruled out in an unconstrained arc by Proposition 1. The control switches between $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$ when ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0, where two unconstrained arcs are connected.
III-B Constrained Arcs in Problem (1a)
Compared to unconstrained arcs, constrained arcs exhibit more complex behavior. In a constrained arc, $\exists p\in\left[P\right]$, s.t. ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$. The following proposition characterizes a constrained arc.
@@ III-B Constrained Arcs in Problem (1a)
Proof.
(2) implies ${\boldsymbol{c}}_{p}^{\top}\left[{\boldsymbol{b}},{\boldsymbol{A}}{\boldsymbol{b}},{\boldsymbol{A}}^{2}{\boldsymbol{b}},\dots,{\boldsymbol{A}}^{n-1}{\boldsymbol{b}}\right]\not={\boldsymbol{0}}$ since ${\boldsymbol{c}}_{p}\not={\boldsymbol{0}}$; hence, $r_{p}$ exists and serves as the minimum non-zero column index of a non-vzero matrix. For $r\in\left[r_{p}-1\right]$, considering the $r$th-order derivative of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$, (11) holds, and ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}u\right)=0$. Hence, (10) holds.
It is allowed that multiple inequality constraints are active in a constrained arc, as proved in Proposition 3.
@@
Proof.
The proof of (12) is omitted since it is elementary [31].
Assume that $\forall r\in\left[n\right]$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$. Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t. $p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$. Therefore, $\forall r\geq n$, $\exists p_{r}^{*}$ s.t. $p_{r}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{r}^{*}\left({\boldsymbol{A}}\right)$.
@@
[…]
@@
Proof.
Assume that $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds in the constrianed arc. By Proposition 2, (10) holds $\forall p\in\mathcal{P}$. $\forall q\in\mathcal{P}$, since ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}\equiv{\mathrm{const}}$, (15) can be derived through considering the $r$th-order derivative of ${\boldsymbol{c}}_{q}^{\top}{\boldsymbol{x}}$ for $r\in\left[n\right]$.
@@
Consider a chain-of-integrator system of order $n$, i.e.,
[…]
Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$. $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent, i.e., a constrained is allowed where both $x_{1}+x_{3}\equiv 0$ and $x_{2}\equiv 0$ hold. In this case, $x_{2}\equiv 0$ implies $x_{1}\equiv 0$ and $u\equiv 0$; hence, $x_{1}+x_{3}\leq 1$ implies $x_{3}\equiv 1$. Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are inconsistent, i.e., in a constrained arc, if $x_{1}+x_{3}\equiv 1$ holds, then $x_{2}\equiv 1$ cannot hold.
III-C System Behavior
@@
For a constrained arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$, $\mathcal{P}\not=\varnothing$. (10) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}$, where $\widehat{{\boldsymbol{a}}}_{p}$ is defined in (10). Although it allows that $\exists p,q\in\mathcal{P}$, s.t. $\widehat{{\boldsymbol{a}}}_{p}\not=\widehat{{\boldsymbol{a}}}_{q}$, it is guaranteed by the consistency of $\mathcal{P}$ that $\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{a}}}_{q}^{\top}{\boldsymbol{x}}$ during the arc. For convenience, let $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$. Then, ${\mathcal{S}}$ in Definition 1 is well-defined. By Lemma 1, during ${\mathcal{S}}$,
@@
Consider a chain-of-integrator system (16) with full box state constraints, i.e., $\forall k\in\left[n\right]$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $-x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. The $\left(2k-1\right)$-th constraint is $-x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$th-order, where $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$, and $u\equiv 0$. Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$ where $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i\in\left[n\right],j\in\left[n\right]}$. $\hat{a}_{ij}=1$ if $k<i=j+1\leq n$; otherwise, $\hat{a}_{ij}=0$.
@@
III-D Switching Law
Based on the analysis in Section III-C, if the optimal trajectory consists of a finite number of arcs, then the optimal trajectory can be represented by a sequence of system behaviors. The switching law is defined as follows:

### v3

@@ Abstract
Most existing necessary conditions for optimal control based on adjoining methods require both state and costate information, yet the unobservability of costates for a given feasible trajectory impedes the determination of optimality in practice. This paper establishes a novel theoretical framework for time-optimal control of controllable linear systems with a single input, proposing the augmented switching law (ASL) that represents the input control and the feasibility in a compact form. Given a feasible trajectory, the perturbed trajectory under the constraints of ASL is guaranteed to be feasible, resulting in a novel state-centric necessary condition without dependence on costate information. A first-order necessary condition is proposed that the Jacobian matrix of the ASL is not of full row rank, which also results in a potential approach to optimizing a given feasible trajectory with the preservation of arc structures. The proposed necessary condition is applied to high-order chain-of-integrator systems with full box constraints, contributing to some theoretical results challenging to reason by costate-based conditions.
@@ I Introduction
Figure 1: Graphical roadmap of the proposed theoretical framework. Main contributions regarding representations and optimal conditions are highlighted in red.
Based on PMP, one can get rid of dependence on costates by deriving the optimal trajectory’s structure. Switching surfaces can be constructed in the state space for time-optimal control of linear systems. Walther et al. [17] computed switching surfaces for some systems using a Gröebner basis. Patil et al. [18] represented switching surfaces using state equations for single-input diagonal linear systems without state constraints. However, for systems with state constraints, it remains challenging to completely construct switching surfaces due to the lack of optimal switching laws. For example, He et al. [19] provided explicit equations for switching surfaces of 3rd-order chain-of-integrator systems with full state constraints, limited to zero terminal states. Wang et al. [20] extended [19] to non-zero terminal conditions, failing to achieve time-optimality for higher-order systems. Numerical methods [21] face the curse of dimensionality. Hence, there is a lack of efficient and general optimality conditions without dependence on costates for high-order systems with state constraints.
Another type of necessary condition leverages Bellman’s principle of optimality [22], asserting that a sub-arc of an optimal trajectory is itself optimal for the induced sub-problem, without explicitly involving costates. With wide applications in optimal control [23, 24], the Hamilton-Jacobi-Bellman (HJB) equation [25] serves as a fundamental necessary condition which is also sufficient under suitable regularity conditions. Evans and James [26] investigated the HJB equation for time-optimal control. Wolenski and Zhuang [27] showed that the minimal time function is the unique proximal solution to the HJB equation. However, most numerical HJB-based methods require high computational cost [28]. From a theoretical perspective, HJB-based methods also fail to preserve arc structures, which hinders the analytical reasoning.
This paper sets out to establish a theoretical framework for time-optimal control of controllable single-input linear systems and develops a novel state-centric necessary condition that requires state information without dependence on costate information. A graphical roadmap is shown in Fig. 1. Sections III and IV derive trajectories’ structures on arcs and keypoints for feasibility, respectively. Guided by the knowledge of trajectory structures, Section V develops the state-centric necessary condition which is applied from theoretical and numerical perspectives in Sections VI and VII as examples, respectively. Proofs of propositions and theorems are provided in Appendix A. The contributions of this paper are as follows:
@@
    Based on the developed framework, this paper proposes a novel state-centric necessary condition for time-optimality, asserting that the Jacobian matrix induced by the ASL is not of full row rank. The condition is derived from the fact that if a BBS trajectory is optimal, then all perturbed feasible trajectories have a strictly longer terminal time than the original one. Compared to existing costate-based conditions, the proposed state-centric necessary condition only requires state information without dependence on costates. Note that costates are usually unobservable in practice, especially when the trajectory is not planned by costate-based optimization.
@@
    This paper proposes a potential optimization idea based on the developed necessary condition. Given a feasible BBS trajectory as the initial value, the motion time is optimized with a fixed ASL until the Jacobian matrix is not of full row rank. During iteration, the feasibility is guaranteed based on the developed ASL framework. Compared to discretized methods, the proposed approach can preserve the BBS arc structure and has the potential to significantly reduce control oscillations and open-loop integration errors of terminal states.
@@
    The proposed state-centric condition is applied to high-order chain-of-integrator systems with full state constraints which remains an open problem in the optimal control domain. An upper bound on the number of arcs is determined in a general sense, while case-by-case analysis is necessary in traditional methods. A recursive equation for the junction time when the state is tangent to the constraints’ boundaries is derived in chattering arcs induced by 2nd-order constraints, which is challenging to prove by costate analysis. As a result, one could determine the existence of chattering induced by 2nd-order constraints in chain-of-integrator systems in future works. The above results demonstrate the theoretical effectiveness of the proposed necessary condition.
@@ II Problem Formulation
$\min_{u}\quad$ ${t_{\mathrm{f}}}=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (1a)
@@
$u\left(t\right)\in\left[-{u_{\mathrm{m}}},{u_{\mathrm{m}}}\right]$ is undetermined during ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$, which is called a singular arc. The notation $\equiv$” means that the equality holds for a continuous period. (7) is the well-known bang-bang and singular (BBS) control law [11, 29]. In other words, the optimal control is always $\pm{u_{\mathrm{m}}}$, except during singular arcs.
Note that the objective function ${t_{\mathrm{f}}}=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t$ is in a Lagrangian form; hence, the continuity of the system is guaranteed by the following equality, i.e.,
@@ III Arc Representation of the Optimal Trajectory
An arc of a trajectory is defined as a continuous segment conditioned by a fixed dynamic equation, i.e., $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}$ where $\widehat{{\boldsymbol{A}}}$ and $\widehat{{\boldsymbol{b}}}$ are constant. Two kinds of arcs, i.e., unconstrained arcs and constrained arcs, are discussed in Section III-A. Through discussion on the two arcs, the system behavior representing a single arc is defined in Section III-B. Finally, an optimal non-chattering trajectory can be represented by a switching law, i.e., a sequence of system behaviors, as discussed in Section III-C.
III-A Unconstrained and Constrained Arcs in Problem (1a)
In an unconstrained arc, ${\boldsymbol{C}}{\boldsymbol{x}}+{\boldsymbol{d}}<{\boldsymbol{0}}$ holds almost everywhere (a.e.), and the control does not switch. Proposition 1 states that $u\equiv\pm{u_{\mathrm{m}}}$ holds in an unconstrained arc.
@@ III-A Unconstrained and Constrained Arcs in Problem (1a)
In an unconstrained arc, $\dot{\boldsymbol{\lambda}}=-{\boldsymbol{A}}^{\top}{\boldsymbol{\lambda}}$ holds since ${\boldsymbol{\eta}}\equiv{\boldsymbol{0}}$. Two cases could occur: (a) ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}\equiv 0$ holds for a period; (b) ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0 for finite times. Case (a) is ruled out in an unconstrained arc by Proposition 1. In Case (b), the control switches between $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$ when ${\boldsymbol{b}}^{\top}{\boldsymbol{\lambda}}$ crosses 0, where two unconstrained arcs are connected.
Compared to unconstrained arcs, constrained arcs exhibit more complex behavior. In a constrained arc, $\exists p\in\left[P\right]$, s.t. ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$. Proposition 2 characterizes a constrained arc. Proposition 3 points out that multiple inequality constraints are allowed to be active in a constrained arc.
@@
Consider a constrained arc for $t\in\left[t_{1},t_{2}\right]$. $\mathcal{P}\subset\left[P\right]$. Assume that during the arc, $\forall p\in\mathcal{P}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds. Then, $\forall p\in\mathcal{P}$, $t\in\left[t_{1},t_{2}\right]$, it holds that
@@
Consider a chain-of-integrator system of order $n$:
$\begin{dcases}\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right],\,1<k\leq n,\\ \dot{x}_{1}\left(t\right)=u\left(t\right),&t\in\left[0,{t_{\mathrm{f}}}\right].\end{dcases}$ (13)
Two inequality constraints are introduced as $x_{1}+x_{3}\leq 1$ and $x_{2}\leq 0$. $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 0$ are consistent, i.e., a constrained arc is allowed where both $x_{1}+x_{3}\equiv 0$ and $x_{2}\equiv 0$ hold. In this case, $x_{2}\equiv 0$ implies $x_{1}\equiv 0$ and $u\equiv 0$; hence, $x_{1}+x_{3}\leq 1$ implies $x_{3}\equiv 1$. Similarly, $x_{1}+x_{3}\equiv 1$ and $x_{2}\equiv 1$ are inconsistent, i.e., in a constrained arc, if $x_{1}+x_{3}\equiv 1$ holds, then $x_{2}\equiv 1$ cannot hold.
III-B System Behavior
@@
For a constrained arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$, $\mathcal{P}\not$$\varnothing$. (9) implies that $\forall p\in\mathcal{P}$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}+{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}$, where $\widehat{{\boldsymbol{a}}}_{p}$ is defined in (9). Although it allows that $\exists p,q\in\mathcal{P}$, s.t. $\widehat{{\boldsymbol{a}}}_{p}\not=\widehat{{\boldsymbol{a}}}_{q}$, it is guaranteed by the consistency of $\mathcal{P}$ that $\widehat{{\boldsymbol{a}}}_{p}^{\top}{\boldsymbol{x}}\equiv\widehat{{\boldsymbol{a}}}_{q}^{\top}{\boldsymbol{x}}$ during the arc. Denote by $\widehat{{\boldsymbol{A}}}\triangleq\widehat{{\boldsymbol{A}}}_{\min\left(\mathcal{P}\right)}$. Then, $\mathcal{S}$} in Definition 1 is well-defined. By Lemma 1, it holds during $\mathcal{S}$} that
@@
Consider a chain-of-integrator system (13) with full box state constraints, i.e., $\forall k\in\left[n\right]$, $t\in\left[0,{t_{\mathrm{f}}}\right]$, $x_{\mathrm{m}k}\leq x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. The $\left(2k-1\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$, and the $\left(2k\right)$-th constraint is $x_{k}\left(t\right)\leq x_{\mathrm{m}k}$. A constrained arc $x_{k}\equiv x_{\mathrm{m}k}$ is of $k$th-order, where $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$, and $u\equiv 0$. Then, $x_{k}\equiv x_{\mathrm{m}k}$ can be represented as ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},{\boldsymbol{0}},{\boldsymbol{I}}_{k\times n},x_{\mathrm{m}k}{\boldsymbol{e}}_{k},\left\{2k\right\}\right)$ where $\widehat{{\boldsymbol{A}}}=\left(\hat{a}_{ij}\right)_{i\in\left[n\right],j\in\left[n\right]}$. $\hat{a}_{ij}=1$ if $k<i=j+1\leq n$; otherwise, $\hat{a}_{ij}=0$.
@@
III-C Switching Law
Based on the analysis in Section III-B, if the optimal trajectory consists of a finite number of arcs, then the optimal trajectory can be represented by a sequence of system behaviors. The switching law is defined as follows: